Another method of solving quadratic eqautions came from India.
a. move the constant term to the right side of the equation.
b. multiply each term in the equation by four times the coefficient of the x square term.
c. Square the coefficient of the orginal x term and add it to both sides.
e. sel the left side of the equation equal to the postive square root of the number on the right side and solve for x.
f. set the left side of the equation equal to the negative square root of the number on the right side of the equation and solve for x.
Solve the following.
x square - 2x -13=0
4x square - 4x + 3 = 0
x square + 12x-64=0
2x square - 3x - 5=0
I need help with any or all of the following to see an example of the process.

Answer by user_dude2008(1862)

x^2 - 2x -13=0
a. x^2 - 2x =13
b. 4x^2-8x=52
c. 4x^2-8x+4=52+4

4x^2-8x+4=56
d. 4(x^2-2x+1)=56
d. 4(x-1)^2=56

(x-1)^2=14

e. x-1=sqrt(14)
e. x=1+sqrt(14)

f. x-1=-sqrt(14)
f. x=1-sqrt(14)


Answers: x=1+sqrt(14) or x=1-sqrt(14)